A rectangular block stands 12 inches tall with a diagonal measuring 13 inches. What is is the length of the base?
Answer by addingup(3677)

A rectangle split diagonally gives you two right triangles.
----------------------------------
Pythagoras:
a^2+b^2= c^2 This is used to find the hypotenuse of the right triangle. Since we want the short side, let's divide both sides by b^2:
a^2= c^2/b^2 Now let's use our numbers:
a^2= 13^2/12^2
a^2= 169/144
a^2= 25 Take the square root on both sides:
a= 5 is the length of the base